# Benchmarking AI Models in Software Engineering: A Review, Search Tool, and Enhancement Protocol

### v1

@@ III-B Results of the Review
Aider  [^9] https://github.com/Aider-AI/aider/blob/main/benchmark/README.md
NL2ML-lib [57]
RMCBench [58]
Evil [59]
Text2Text (about code) InfiCoder-Eval [60]
BRIGHT [43]
Code2Text DeepCom [61]
Hybrid-DeepCom [62]
BinSum [63]
Code Attention [64]
Funcom [65]
CodeSum [66]
CoDesc [67]
Parallel [68]
CoDocBench [69]
TABLE VII: Overview of SQL-related Benchmarks.
Category Name
Text2Code (SQL) BIRD [70]
KaggleDBQA [71]
StacQc [72]
Spider(V2) [73]
Spider-Syn [74]
Spider-Real [75]
Spider-DK [76]
Spider-CN [77]
SParC [78]
Lyra [79]
DuSQL [80]
CoSQL [81]
@@
Table XX features benchmarks with competitive programming as their root, i.e., those used for understanding code complexity and efficiency. Table XXI features a set of benchmarks specifically designed to evaluate the performance of models on Data Science-related tasks along with some other domain-specific SE tasks. A notable example in this table is the Bio-Coder series of benchmarks specifically designed for bioinformatics tasks. To assess the mathematical reasoning capabilities of AI4SE models, see Table XXII. Besides numbers and code, natural language is also a key component in AI4SE. From supporting instruction-tuned AI4SE models, which align more with the human brain [82], that aim to accomplish question and answering (QA) similar to the widely recognized platform StackOverflow, to summarizing code and generating tags, Table XXIII features AI4SE benchmarks including natural language: text-to-code, code-to-text and text-to-text (code related). We have intentionally isolated all SQL-related benchmarks due to their abundance and to facilitate locating the correct benchmark; these are presented in Table XXIV.
TABLE VIII: Overview of Programming Language Translation Benchmarks.
Category Name
Programming Languages CodeTrans [49]
TransCoder-ST [83]
CoST [84]
AVATAR [85]
Multilingual-Trans [86]
NicheTrans [86]
LLMTrans [86]
G-TransEva [87]
CODEDITOR [88]
Libraries DLTrans [86]
Intermediate Representation SLTrans [89]
Language Conversion Frameworks MultiPL-E [6]
MultiEval [5]
@@
TABLE IX: Overview of Selected Real-to-Life SE Benchmarks.
Category Benchmark
@@
Software Development & Agent Benchmarks DevBench [90]
DevEval [91]
CoderUJB [92]
CODAL [25]
ToolQA [93]
MIT [94]
SAFIM [95]
AgentBench [96]
Class Level ClassEval [97]
CONCODE [50]
BigCodeBench [98]
Project & Cross-file SWE-bench [99]
CrossCodeEval [100]
CoderEval [101]
DotPrompts [102]
BigCloneBench [103]
DI-Bench [104]
DyPyBench [105]
Repository Level RepoBench [106]
RepoEval [107]
EvoCodeBench [108]
SketchEval [109]
Stack Repo [109]
ML-BENCH [110]
CodeGen4Libs [111]
TABLE X: Overview of Selected API and Retrieval Benchmarks by Category.
Category Benchmark
API Prediction RestBench [112]
APIBench-Q [113]
BIKER [114]
Gorilla APIBench [115]
Gorilla APIZoo [115]
Retrieval & Planning API-Bank [116]
CodeRAG-Bench [117]
Search4Code [118]
CoIR [119]
Memorization SATML-ext [120]
TABLE XI: Overview of AI4SE Benchmarks Related to Pseudocode.
Category Benchmark
Pseudocode to Code SPoC [121]
NAPS [122]
Code to Pseudocode Django [123]
TABLE XII: Overview of Selected Crowd-sourced Benchmarks.
Category Benchmark
Crowd-sourced Benchmarks WikiSQL [124]
Spider [125]
NL2Bash [126]
NAPS [122]
SPoC [121]
MBPP [14]
@@
While translating natural language is more trivial nowadays, translating code remains challenging due to various reasons (e.g. versioning, semantics, dependencies). With the lack of diversity in language support for AI4SE benchmarks and also benefiting numerous other SE tasks, Table XXV features an overview of resources that can support the ongoing development of code translation.
@@
However, the workflow of a developer is not merely an exercise of writing snippets of code for given descriptions, but rather having a general overview of a project and how one can implement functionality such that it fits well into a collective code base. For this, Table XXVI provides a collection of benchmarks that examine the capabilities of models to generate code on a larger scale.
@@
The utilization of APIs plays a significant role in AI4SE benchmarks, specifically for models with Retrieval Augmented Generation (RAG) capabilities. In Table XXVII, prominent benchmarks focusing on leveraging the power of APIs are denoted. Furthermore, Table XXVIII lists benchmarks related to pseudocode, followed by an overview of notable crowd-sourced AI4SE resources in Table XXIX.
@@
TABLE XIII: Overview of Automated Program Repair, Fault Localization, and Vulnerability Detection Benchmarks.
Category Benchmark
@@
Automated Program Repair & Fault Localization Defects4J [127]
GitBug-Java [128]
EvalGPTFix  [129]
TutorCode  [130]
GHRB  [131]
IntroClass  [132]
ManyBugs  [132]
DebugBench [133]
QuixBugs [134]
RES-Q [135]
StudentEval [136]
Re-Factory [137]
ConDefects [138]
Cerberus [139]
Vulnerability Detection CVEFixes [140]
LLMSecEval [141]
SecurityEval [142]
Vul4J [143]
FormAI [144]
VJBbench [145]
SmartBugs [146]
Devign [147]
D2A [148]
BigVul [149]
SARD  [^10] https://samate.nist.gov/SARD/
Juliet 1.3  [^11] https://samate.nist.gov/SARD/test-suites/112
NVD  [^12] https://nvd.nist.gov/developers/data-sources
Software Testing CoverageEval [150]
ATLAS [151]
HITS [152]
MeMo [153]
MLAPIs [154]
TABLE XIV: Overview of Selected SE-Workflow Benchmarks.
Category Benchmark
Code Synthesis & Understanding Methods2Test [155]
CRUXEval [156]
CRQBench [157]
CriticBench [158]
CodeScope [159]
Merge Conflict Repair ConflictBench [160]
Type Inference TypeEvalPy  [161]
TypeEvalPy AutoGen  [161]
@@
Automatic Code Quality Review CodeReview  [162]
Software Maintainability  [163]
Hallucination Detection HALLUCODE [164]
@@
With AI4SE models mainly being utilized for program synthesis, it remains relatively questionable how effective these models are in generating tests and repairing bugs, as it is unclear whether these models truly understand code. For example, Siddiq et al. [165] observed Codex [1] being able to get above 80% coverage for HumanEval [1], yet many test smells were discovered and for another dataset, no higher than 2% coverage was attained. This reveals the importance of benchmarking AI4SE models’ capabilities in test generation, bug repair, and understanding. In Table XXX, several benchmarks are listed that make an effort to assess the aforementioned. Additionally, Table XXXI table features benchmarks that have been designed to evaluate a model’s capabilities in dealing with everyday tasks of a software engineer (e.g., merge-conflict repair, Code Reviews, etc.).
@@
TABLE XV: Overview of Multi-Category Benchmarks, Covering Various Tasks.
Category Name
Multi-Task Benchmark Big-Bench [166]
Cross-Language Code Tasks XLCoST [167]
Cross-Language Code Benchmark CrossCodeBench [168]
Long Code Tasks Long Code Arena [169]
Code Search & Documentation CodeXGLUE [169]
MicrosoftDocs [^13] https://github.com/MicrosoftDocs/
CodeSearchNet [170]
@@
Finally, we collect a set of generic benchmarks in Table XXXII. These benchmarks are not only single-task benchmarks like others previously seen in this section but are rather a collection of tasks spanning a wide range of languages and task types. A notable, and widely known benchmark in this category is the BigBench Benchmark [166] which consists of 167 tasks (not all relevant to AI4SE).

### v2

@@ 3.2. Results of the Review
Aider  [^10] https://github.com/Aider-AI/aider/blob/main/benchmark/README.md English
$\rightarrow$ {C++, GO, Java, 225 problmes
JS, Python, Rust}
NL2ML-lib (Shin et al., 2024) English $\rightarrow$ Python (ML libraries) 11,000
RMCBench (Chen et al., 2024) English $\rightarrow$ 9 Languages 473 malicious prompts
Evil (Liguori et al., 2021) English $\rightarrow$ {Python, IA_32} 19255
Exec-CSN (Xie et al., 2024) English $\rightarrow$ Python 1,931
CodeIF (Yan et al., 2025) English $\rightarrow$ {Java, Python, Go, C++} 1,200 tasks
CodeIF-Bench (Wang et al., 2025e) English $\rightarrow$ Python 122 tasks (42 repositories)
ARCADE (Yin et al., 2023) English $\rightarrow$ Python 1082
StackEval (Shah et al., 2024) English $\rightarrow$ Python 925
SwiftEval (Petrukha et al., 2025) English $\rightarrow$ Swift 28
CoSQA+ (Gong et al., 2025) English $\rightarrow$ Python 412,080 pairs
Table 7. Overview of Natural Language Benchmarks (Continued).
Category Name Language(s) No. of Problems
Text2Text (about code) InfiCoder-Eval (InfiCoder, 2023) English $\rightarrow$ English 270
BRIGHT (Su et al., 2024) English $\rightarrow$ English 1,398
Code2Text DeepCom (Hu et al., 2018) Java $\rightarrow$ English 588K
Hybrid-DeepCom (Hu et al., 2020) Java $\rightarrow$ English 466k
BinSum (Jin et al., 2023) Binary functions $\rightarrow$ English 557K
Code Attention  (Allamanis et al., 2016) Java $\rightarrow$ English 11 projects
Funcom  (LeClair et al., 2019) Java $\rightarrow$ English 2.1M problems
CodeSum (Hu et al., [n. d.]) Java $\rightarrow$ English 410,630
CoDesc (Hasan et al., 2021) Java $\rightarrow$ English 4.21M datapoints
Parallel (Barone and Sennrich, [n. d.]) Python $\rightarrow$ English 150k function/doc pais
CoDocBench (Pai et al., 2025) Python $\rightarrow$ English 4573 code/doc pairs
PoorCodeSum (Hu et al., 2024) {Java, Python, Go} {10,955 , 14,918 , 8,122}
$\rightarrow$ English
P-CodeSum (Yun et al., 2024) {Python, Java, Go, 1,500 pairs
JS, PHP, Ruby} $\rightarrow$ English
Table 8. Overview of SQL-related Benchmarks.
Category Name Language(s) No. of Problems
Text2Code BIRD  (Li et al., 2023b) English $\rightarrow$ SQL 12,751
KaggleDBQA  (Lee et al., 2021) English $\rightarrow$ SQL 272, paired with golden solutions
StacQc  (Yao et al., 2018) English $\rightarrow$ {Python/SQL} {147,546 / 119,519}
question-answer pairs
Spider(V2111111see Table 13(Lei et al., 2024) English $\rightarrow$ SQL 632 queries
Spider-Syn (Gan et al., 2021b) English $\rightarrow$ SQL (7000 / 1034)
Spider-Real (Deng et al., 2021) English $\rightarrow$ SQL 508
Spider-DK (Gan et al., 2021a) English $\rightarrow$ SQL 535 pairs
Spider-CN (Min et al., 2019) Chinese $\rightarrow$ SQL 9691 queries
SParC (Yu et al., 2019b) English $\rightarrow$ SQL 4,298 question sequences
Lyra (Liang et al., 2022) {English, Chinese} $\rightarrow$ {python, SQL} 2000
DuSQL (Wang et al., 2020) Chinese $\rightarrow$ SQL 23,797 question/SQL pairs
CoSQL (Yu et al., 2019a) English $\rightarrow$ SQL 3,007 Question Sequences
SynSQL-2.5M (Li et al., 2025a) English $\rightarrow$ SQL 2,544,390
PAUQ (Bakshandaeva et al., 2022) Russian $\rightarrow$ SQL 9,691
Ar-Spider (Almohaimeed et al., 2024) Arabic $\rightarrow$ SQL 9,691
Tur2SQL (Kanburoğlu and Tek, 2023) Turkish $\rightarrow$ SQL 10,809 question/SQL pairs
@@
Table 3 features benchmarks with competitive programming as their root, i.e., those used for understanding code complexity and efficiency. Table 4 features a set of benchmarks specifically designed to evaluate the performance of models on Data Science-related tasks along with some other domain-specific SE tasks. A notable example in this table is the Bio-Coder series of benchmarks specifically designed for bioinformatics tasks. To assess the mathematical reasoning capabilities of AI4SE models, see Table 5. Besides numbers and code, natural language is also a key component in AI4SE. From supporting instruction-tuned AI4SE models, which align more with the human brain (Aw et al., 2023), that aim to accomplish question and answering (QA) similar to the widely recognized platform StackOverflow, to summarizing code and generating tags, Table 6 and Table 7 feature AI4SE benchmarks including natural language: text-to-code, code-to-text and text-to-text (code related). We have intentionally isolated all SQL-related benchmarks due to their abundance and to facilitate locating the correct benchmark; these are presented in Table 8.
@@
Table 9. Overview of Selected Real-to-Life SE Benchmarks. (Note: X/Y/Z denotes Train/Dev/Test)
Category Benchmark Language(s) No. of Problems
@@
Software Development & Agent Benchmarks DevBench (Li et al., 2024e) Python, C/C++, Java, JavaScript 22 repositories
DevEval (Li et al., 2024d) Python 1,874
CoderUJB (Zeng et al., 2024a) Java 2,239
CODAL (Weyssow et al., 2024) Python 500
ToolQA (Zhuang et al., 2023) Python, Math, English 800(Easy)/730(Hard)
MIT (Wang et al., 2024d) Python, English 586 Problems
SAFIM (Gong et al., 2024) Python, Java, C++, C# 17720
AgentBench (Liu et al., 2023d) N/A 1360 prompts
CSR-Bench (Xiao et al., 2025) Python 100 repositories
Class Level ClassEval (Du et al., 2023) Python 100
CONCODE (Iyer et al., 2018) English, Java 104,000
BigCodeBench (Zhuo et al., 2024) Python 1,140
OOP-Bench (Wang et al., 2024b) Python 431
CodeSense (Roy et al., 2025) Python, C, Java 2125 (Python), 876 (C)
875 (Java
ClassEval-T (Xue et al., 2025) Java, C++ 1,243)
Project & Cross-file SWE-bench (Jimenez et al., 2023) Python 19,008 (Train), 225 (Dev),
2,294 (Test)
144 (Small)
CrossCodeEval (Ding et al., 2023) C#, TypeScript, Java, Python 2,665 (Python), 2,139 (Java),
3,356 (TypeScript), 1,768 (C#)
CoderEval (Yu et al., 2023) Java, Python 230
DotPrompts (Agrawal et al., 2023) Java 105538 problems
(1420 methods)
BigCloneBench (Svajlenko and Roy, 2015) Java 25,000 Java Systms
DI-Bench (Zhang et al., 2025c) Python, C#, Rust, JS 581 repositories
(w/ dependencies)
DyPyBench (Bouzenia et al., 2024) Python 50 repositories
Multi-SWE Java, TypeScript, JavaScript 500 (Python), 128 (Java),
-bench (Zan et al., 2025) Go, Rust, C, C++ 224 (TS), 356 (JS),
428 (Go), 239 (Rust)
128 (C), 129 (C++)
KernelBench (Ouyang et al., 2025) Python 250
CodeMEnv (Cheng et al., 2025) Python, Java 587 (Python), 335 (Java)
CodeEditorBench (Guo et al., 2025) Python, Java, C++ 7,961
ProjectEval (Liu et al., 2025a) Python 284
Table 10. Overview of Selected Real-to-Life SE Benchmarks. (Continued - Note: X/Y/Z denotes Train/Dev/Test)
Category Benchmark Language(s) No. of Problems
Repository Level RepoBench (Liu et al., 2023b) Python, Java Cross-file: 8,033, In-file: 7,910
RepoEval (Zhang et al., 2023a) Python 1,600 (line),
1,600 (API), 373 (function)
EvoCodeBench (Li et al., 2024c) Python 275
SketchEval (Zan et al., 2024) Python 19 repositories
(5 easy, 8 medium, 6 hard)
Stack Repo (Zan et al., 2024) Python (435,890 / 220,615 / 159,822)
answer pairs
ML-BENCH (Tang et al., 2024a) Python & Bash 9641 problems
CodeGen4Libs (Liu et al., 2023c) Java 403,780 prompts
SWE-rebench (Badertdinov et al., 2025) Python 294 tasks, 169 repositories
SWE-Polybench (Rashid et al., 2025) Java, JavaScript 165 (Java), 1017 (JavaScript),
TypeScript, Python 729 (TypeScript), 199 (Python)
HumanEvo (Zheng et al., 2025b) Python, Java 200 (Python), 200 (Java)
REPOCOD (Liang et al., 2024) Python 980 functions (11 projects)
FEA-Bench (Li et al., 2025b) Python 83 repositories
JavaBench (Cao et al., 2024) Java 389 methods (106 classes)
SWE-bench live (Zhang et al., 2025b) Java 1,319 tasks (93 repositories)
SWE-Lancer (Miserendino et al., 2025) N/A 1,488 tasks)
Table 11. Overview of Selected API and Retrieval Benchmarks by Category.
Category Benchmark Sources/API(s) No. of Problems
API Prediction RestBench (Song et al., 2023) Spotify, TMDB 57, 100
APIBench-Q (Peng et al., 2021) StackOverflow, Tutorial Websites 6,563 (Java),
4,309 (Python)
BIKER (Huang et al., 2018) StackOverflow 33,000
Gorilla APIBench (Patil et al., 2023) HuggingFace, TensorHub, TorchHub 925, 696, 94
Gorilla APIZoo (Patil et al., 2023) Open submissions
(Google, YouTube, Zoom, etc.)
Retrieval & Planning API-Bank (Li et al., 2023d) 73 commonly used APIs 753
CodeRAG-Bench (Wang et al., 2024a) Competition solutions, tutorials, 25,859
documentation, StackOverflow, GitHub
Search4Code (Rao et al., 2021) Bing 6596(java)/4974(c#)
CoIR (Li et al., 2024a) GitHub, StackOverflow, and 2.38M (corpus)
Various Benchmarks 3.37(queries)
Memorization SATML-ext (Al-Kaswan et al., 2024) GitHub 1,000 samples
API Misuse Detection ExampleCheck (Zhang et al., 2018) StackOverflow 100 (Java)
ROBUSTAPI (Zhong and Wang, 2024) StackOverflow 1208 (18 Java APIs)
APIMU4C (Gu et al., 2019) Juliet Test Suite, ITC, 2272 (C)
OpenSSL, Curl, Httpd
Table 12. Overview of AI4SE Benchmarks Related to Pseudocode.
Category Benchmark Language(s) No. of Problems Crowdsourced
Pseudocode to Code SPoC (Kulal et al., 2019) C++ 18,356 Yes
NAPS (Zavershynskyi et al., 2018) Java/UAST 17,477 No
PseudoEval (Wu et al., 2025) Python, C++, Rust 1,060 No
Code to Pseudocode Django (Oda et al., 2015) Python, English 18,805 (Train), 1,000 (Dev), No
& Japanese 1,805 (Test)
Table 13. Overview of Selected Crowd-sourced Benchmarks.
Benchmark Language(s) No. of Problems Source
WikiSQL (Zhong et al., 2017) Natural language $\rightarrow$ SQL query 80,654 Amazon MTurk (deprecated - 2017)
Spider (Yu et al., 2018) Natural language $\rightarrow$ SQL query 10,181 11 Yale students (2018)
NL2Bash (Lin et al., 2018) Natural language $\rightarrow$ Bash 9,305 Upwork (2018)
NAPS (Zavershynskyi et al., 2018) Java/UAST $\rightarrow$ Pseudocode 17,477 Self-hosted crowdsourcing,
comp. programming community (2018)
SPoC (Kulal et al., 2019) C++ 18,356 comp. programming websites (2019)
MBPP (Austin et al., 2021) Python 974 Google Research,
internal crowdworkers (2021)
@@
While translating natural language is more trivial nowadays, translating code remains challenging due to various reasons (e.g. versioning, semantics, dependencies). With the lack of diversity in language support for AI4SE benchmarks and also benefiting numerous other SE tasks, Table 14 features an overview of resources that can support the ongoing development of code translation.
Table 14. Overview of Programming Language Translation Benchmarks (Note: X/Y/Z denotes Train/Dev/Test).
Category Name Language(s) No. of Samples
Programming Languages CodeTrans (Lu et al., 2021) C#, Java 11,800
TransCoder-ST (Roziere et al., 2022) C++, Java, Python 437,030
CoST (Zhu et al., 2022b) 7 programming languages 16,738
AVATAR (Ahmad et al., 2023) Java, Python 7,133 / 476 / 1,906
Multilingual-Trans (Yan et al., 2023) 8 programming languages 30,419 total
NicheTrans (Yan et al., 2023) Various niche languages 236,468 total
LLMTrans (Yan et al., 2023) 8 programming languages 350
G-TransEva (Jiao et al., 2023) 5 programming languages 400 total
CODEDITOR (Zhang et al., 2023b) C# & Java 6613
RustRepoTrans (Ou et al., 2024) C++, Java, Python $\rightarrow$ Rust 375
AVATAR-TC (Jana et al., 2024) Java, Python 55,179 / 443 / 1,746
RustRepoTrans (Ou et al., 2024) C++, Java, Python $\rightarrow$ Rust 375
RepoTransBench (Wang et al., 2024e) Python $\rightarrow$ Java 100 repositories
Libraries DLTrans (Yan et al., 2023) PyTorch, TensorFlow,
MXNet, Paddle 408 total
Intermediate Representation SLTrans (Paul et al., 2024) 14 Languages $\rightarrow$ LLVM-IR 4M
Language Conversion Frameworks MultiPL-E (Cassano et al., 2022) 19 programming languages -
MultiEval (Athiwaratkun et al., 2022) 13 programming languages -
@@
However, the workflow of a developer is not merely an exercise of writing snippets of code for given descriptions, but rather having a general overview of a project and how one can implement functionality such that it fits well into a collective code base. For this, Table 9 and Table 10 provide a collection of benchmarks that examine the capabilities of models to generate code on a larger scale.
@@
The utilization of APIs plays a significant role in AI4SE benchmarks, specifically for models with Retrieval Augmented Generation (RAG) capabilities. In Table 11, prominent benchmarks focusing on leveraging the power of APIs are denoted. Furthermore, Table 12 lists benchmarks related to pseudocode, followed by an overview of notable crowd-sourced AI4SE resources in Table 13.
@@
Table 15. Overview of Automated Program Repair, Fault Localization, and Vulnerability Detection Benchmarks.
Category Benchmark Language(s) No. of Samples
@@
Automated Program Repair & Fault Localization Defects4J (Just et al., 2014) Java 835
GitBug-Java (Silva et al., 2024) Java 199
EvalGPTFix  (Zhang et al., 2023c) Java 4530
TutorCode  (Yang et al., 2024) C++ 1239
GHRB  (Lee et al., 2023) Java 107
IntroClass  (Goues et al., 2015) C 998
ManyBugs  (Goues et al., 2015) 7 Languages 185
DebugBench (Tian et al., 2024) C++, Java, Python 1,438 & 1,401 & 1,414
QuixBugs (Lin et al., 2017) Java 40 (locations of bugs)
RES-Q (Labash et al., 2024) Python, JS 100 hand-crafted questions
+ test scripts
StudentEval (Babe et al., 2023) Python 1,749 buggy programs (48 tasks)
(3 test cases per problem)
Re-Factory (Hu et al., 2019) Python 1783(buggy)/2442(correct)
ConDefects (Wu et al., 2023b) Python, Java 526(Python), Java(477)
Cerberus (Shariffdeen et al., 2023) C, C++, Java 2242 (across 4 task types)
RepairBench (Silva and Monperrus, 2025) Java 574
MaRV (Nunes et al., 2025) Java 693
BugsinPy (Widyasari et al., 2020) Python 493 bugs
Vulnerability Detection CVEFixes (Bhandari et al., 2021) Various Languages 5,365
LLMSecEval (Tony et al., 2023) C 150 (on 25 distinct vulnerabilities)
SecurityEval (Siddiq et al., 2022) 6 languages 130 (on 75 common vulnerabilities)
Vul4J (Bui et al., 2022) Java 79 vulnerabilities
FormAI (Tihanyi et al., 2023) C 112,000 labeled instances
VJBbench (Wu et al., 2023a) Java 42 vulnerabilities
SmartBugs (Durieux et al., 2020) Solidity 69 Vulnerable Smart Contracts
Devign (Zhou et al., 2019) C 4 large-scale software repositories
D2A (Zheng et al., 2021) C/C++ 6 OSS Programs
BigVul (Fan et al., 2020) C/C++ 348 Projects
SARD  [^12] https://samate.nist.gov/SARD/ Java, C, C++, C#, PHP 32k [^13] As of 4th Feb 2025
Juliet 1.3  [^14] https://samate.nist.gov/SARD/test-suites/112 C/C++ 64k [^15] As of 4th Feb 2025
NVD  [^16] https://nvd.nist.gov/developers/data-sources Various Languages 265k [^17] As of 4th Feb 2025
ARVO (Mei et al., 2024) C, C++ 1,001 vulnerabilities
VADER (Liu et al., 2025b) 15 languages 174 vulnerabilities
ManyVuls4J (Lin et al., 2024b) Java 103 vulnerabilities
Software Testing CoverageEval (Tufano et al., 2023) Python 1160
ATLAS (Watson et al., 2020) Java 9,275 projects
HITS (Wang et al., 2024c) Java 10 projects
MeMo (Bareiß et al., 2022) Java 9 projects
MLAPIs (Wan et al., 2022) Python 63 applications
CoderUJB (Zeng et al., 2024b) Java 2,239
TestBench (Zhang et al., 2024b) Java 108
TestEval (Wang et al., 2024f) Python 210
TARBENCH (Yaraghi et al., 2025) Java 45,373 (59 projects)
ProjectTest (Wang et al., 2025c) Python, Java, Javascript 20 per language
CLOVER (Xu et al., 2025c) Python 845
Table 16. Overview of Selected SE-Workflow Benchmarks.
Category Benchmark Language(s) No. of Samples
Code Synthesis & Understanding Methods2Test (Tufano et al., 2020) Java 780,944
CRUXEval (Gu et al., 2024) Python 800
CRQBench (Dinella et al., 2024) C++ 100
CriticBench (Lin et al., 2024a) Python 3,825(across 5 tasks)
CodeScope (Yan et al., 2024) 8 PLs 13,390 (across 8 tasks)
CodeCriticBench (Zhang et al., 2025a) Various Languages 1,517 (Easy), 1,084 (Medium),
1,699 (Hard)
CRUXEval-X (Xu et al., 2025a) 19 PLs 19K
Merge Conflict Repair ConflictBench (Shen and Meng, 2024) Java 180
Type Inference TypeEvalPy  (Venkatesh et al., 2023) Python 845 (annotated labels)
TypeEvalPy AutoGen  (Venkatesh et al., 2023) Python 78373 (annotated labels)
@@
Automatic Code Quality Review CodeReview  (Li et al., 2022b) 8 languages 7.9M pull requests
Software Maintainability  (Schnappinger et al., 2020) Java 519 projects
(evaluations of quality)
BenMark (Wang et al., 2025d) Java 1,299,186 methods
CodeReviewer (Li et al., 2022c) 9 languages 13,100
CodeReview-New (Guo et al., 2024) 16 languages 14,600
CodeReviewQA (Lin et al., 2025) 9 languages 900
Hallucination Detection HALLUCODE (Liu et al., 2024) Python 5,663
CodeHaluEval (Tian et al., 2025) Python 699
Collu-Bench (Jiang et al., 2024a) N/A 13,234
LMDefects (Fan et al., 2023a) Java 60 (Easy), 53 (Medium)
Codemirage (Agarwal et al., 2024) Python 1,137
@@
With AI4SE models mainly being utilized for program synthesis, it remains relatively questionable how effective these models are in generating tests and repairing bugs, as it is unclear whether these models truly understand code. For example, Siddiq et al. (Siddiq et al., 2024) observed Codex (Chen et al., 2021) being able to get above 80% coverage for HumanEval (Chen et al., 2021), yet many test smells were discovered and for another dataset, no higher than 2% coverage was attained. This reveals the importance of benchmarking AI4SE models’ capabilities in test generation, bug repair, and understanding. In Table 15, several benchmarks are listed that make an effort to assess the aforementioned. Additionally, Table 16 table features benchmarks that have been designed to evaluate a model’s capabilities in dealing with everyday tasks of a software engineer (e.g., merge-conflict repair, Code Reviews, etc.).
@@
Table 17. Overview of Multi-Category Benchmarks, Covering Various Tasks.
Name Language(s) Tasks Information
Big-Bench (Srivastava et al., 2022) Python, Numeric, JSON, English Functions over numbers, Mathematical Reasoning, Text2Code, Code2Text, Code explanation, Debugging, Turing Complete Concept Learning, amongst other tasks 250, several per category, 42, 60, 66, 34, 6,390
XLCoST (Zhu et al., 2022a) C, C++, C#, Java, JavaScript, Kotlin, PHP, Python, Ruby, Rust Text2Code (program synthesis, code search), Code Summarization, Code Translation 567K (509k, 58k), 567K, 122K
CrossCodeBench (Niu et al., 2023) Java, C#, Python, C++, JS, PHP, Go, Ruby, TS, C, Bash, Shell Classification, In-Filling, Translation, Generation, Summarization, Type Prediction, Question Answering 6.6M, 13.4M, 2.4M, 19.5M, 11.2M, 773K, 190K
Long Code Arena (Bogomolov et al., 2024) English, Python, Java, Kotlin Commit Message Generation, Module Summarization, Library-Based Code Generation, Project-Level Code Completion, Bug Localization, CI Builds Repair 163, 216, 150, 908 (varying sizes), 14.96K, 78
CodeXGLUE  (Bogomolov et al., 2024) MicrosoftDocs [^18] https://github.com/MicrosoftDocs/ CodeSearchNet (Husain et al., 2019) English, Chinese, Norwegian, Danish, Latvian Go, Java, JavaScript, PHP, Python, Ruby Code Documentation Translation, Code Documentation (Code Summarization, Comment Generation) (CN: 52K, NOR: 26K, DK: 45K, LT: 21K), 621870
DomainEval (Zhu et al., 2025) Python Computation, Network, Basic operation, System, Visualization, Cryptography 5,892 cases total
CodeApex (Liu et al., 2022) C++, English, Chinese Programming comprehension Code generation, Code correction 250, 476, 1330
@@
Additionally, wee collect a set of generic benchmarks in Table 17. These benchmarks are not only single-task benchmarks like others previously seen in this section but are rather a collection of tasks spanning a wide range of languages and task types. A notable, and widely known benchmark in this category is the BigBench Benchmark (Srivastava et al., 2022) which consists of 167 tasks (not all relevant to AI4SE).
Table 18. Overview of Selected Benchmarks for Log Statement Generation and Parsing.
Category Benchmark Language(s) No. of Problems
Log Statement Generation LANCE (Mastropaolo et al., 2022) Java 76,421
LogBench (Li et al., 2023c) Java 6,849
SCLogger (Li et al., 2024b) Java 31,170
AL-Bench (Tan et al., 2025) Java 39,600
Log Parsing LogBase (Zhang et al., 2025d) Various Languages 85,300
LogHub (Zhu et al., 2023) Various Languages 32,000
LogHub-2.0 (Jiang et al., 2024b) Various Languages 3.6M
LogPM (Hashemi et al., 2024) Various Languages 10,821,589
LogEval (Cui et al., 2024) English, Chinese 4,000 log entries across 4 tasks
Finally, we have identified a rather large count of papers that have to do with the matter of Log Parsing and Log Statement generation. we have presented these benchmarks in Table 18.
3.3. Limitations of Existing Benchmarks
Table 19. Common limitations in AI4SE benchmarks
Limitation Description Representative Examples
Language Specific Focuses only on one programming language HumanEval, MBPP, APPS
No Leaderboard No official leaderboard; hinders fair and consistent comparisons SPoC, RepoEval, CONCODE
Poor Maintenance / Data Quality Dataset contains outdated problems, lacks documentation, or is not updated regularly ToolQA, ExampleCheck
Not Peer Reviewed Not been published in a peer-reviewed venue CodeSearchNet, Methods2Test
Infrequently Used Rarely cited in academic work JavaBench, CoIR
Dataset Creation Method How problems were obtained: mined from real-world sources, created by humans, or LLM-generated AVATAR (mined), Spider (created), CRUXEval (generated)
Across our review, we observed several trends in the existing AI4SE benchmarks. The majority are language-specific, most commonly targeting Python or Java. Although these languages are popular in both education and industry, this narrow scope potentially limits their generalization across domains and programming environments. Many benchmarks suffer from poor maintenance, with little to no updates or active support following their initial release. Official and dynamic leaderboards are often missing, which makes it difficult to fairly compare model performance across time or publications, particularly as newer models continue to emerge. Notably, benchmark popularity does not appear to strongly correlate with its maintenance quality or whether it was peer reviewed, suggesting other factors such as visibility or convenience may drive adoption.
There is also variation in how benchmarks are constructed. Manually created datasets are often significantly smaller, whereas newer benchmarks increasingly leverage LLMs to generate large volumes of tasks. Mined datasets which are mostly sourced from GitHub, can be large but are frequently under-specified, with limited transparency into the repositories used or the sampling heuristics applied.
Table 19 summarizes these limitations and provides representative benchmarks for each. The full list of evaluated benchmarks and their metadata can be found in the replication package of the study.